OpenStudy (anonymous):

Evaluate the integral: \[\int\limits_{}^{}ye ^{0.2y}dy\] The problem we are having is the antiderivative of \[e ^{0.2y}.\] We decided to split the problem into: u = y \[uv = e^{0.2y}\] We just aren't sure where to go from here. Help?

OpenStudy (anonymous):

Or rather, what is the integral of (ye^0.2y)dy ?

OpenStudy (anonymous):

I like looking at 0.2 as y/5 it is easier to see how it works... so for this problem try: dv = e^y/5 v = 5e^y/5 u = y du = dy Go for it again!! :D

OpenStudy (anonymous):

Thanks for the help setting it up, again.

OpenStudy (anonymous):

always a pleasure!

OpenStudy (anonymous):

We ended up getting 5ye-25e^y/5dy. Correct? Thanks yet again for the help!

OpenStudy (anonymous):

There shouldn't be a dy term in your answer.. umm, let me see what i get and i'll post it and we can find where you got off :) you're definitely close though!

OpenStudy (anonymous):

sorry it wasn't dy it was 5ye-25e^y/5+C

OpenStudy (anonymous):

oh nvmd i misread what you wrote, it's e^y/5 nothing divided by 5 xD i'm sorry.. everything looks good, just put a ^y/5 power on the first e in your answer :)